# Gordian split links in the Gehring ropelength problem

### v1

@@ 6.2 Circle decompositions - the details
Let $C:I\to\mathbb{R}^{3}$ be a uniformly continuous curve and $\varepsilon>0$. Then there is a $\delta>0$ such that for all $\tilde{C}$ with $||\tilde{C}-C||<\delta$ and all intervals $a,b,a^{\prime},b^{\prime}\in I$ with $a<b$, $a^{\prime}<b^{\prime}$ and $|a-a^{\prime}|,|b-b^{\prime}|<\delta$, we have $||\bar{C}-\bar{C}^{\prime}||<\varepsilon,$ where
@@
$\bar{C}(t)=\begin{cases}C(t)\text{ if }t\in[a,b]\\ C(a)\text{ if }t<a\\ C(b)\text{ if }t>b\end{cases},\quad\bar{C}^{\prime}(t)=\begin{cases}C^{\prime}(t)\text{ if }t\in[a^{\prime},b^{\prime}]\\ C^{\prime}(a^{\prime})\text{ if }t<a^{\prime}\\ C^{\prime}(b^{\prime})\text{ if }t>b^{\prime}\end{cases}.$
@@
Let $C:I\to\mathbb{R}^{3}$ be a uniformly continuous curve and $\varepsilon>0$. Then there exists a $\delta>0$ such that for every $C^{\prime}$ with $||C^{\prime}-C||<\delta$ and every collection of $n$ subcurves $C_{i}:[a_{i},b_{i}]\to\mathbb{R}^{3}$, $i=1,\dots,n$ of $C$ and of $n$ subcurves $C^{\prime}_{i}:[a^{\prime}_{i},b^{\prime}_{i}]\to\mathbb{R}^{3}$ of $C^{\prime}$ with $|a_{i}-a^{\prime}_{i}|,|b_{i}-b^{\prime}_{i}|<\varepsilon$ we have
$||C-C^{\prime}||<\varepsilon,$
@@
Let $\tilde{L}=\tilde{C}_{1}\cup\tilde{C}_{2}$ be a thick link with $||L-\tilde{L}||=d$.
[…]
@@
$||C_{1}^{j}-\tilde{C}_{1}^{j}||<\delta_{2}$
as long as $\mathbf{\delta}$ and $\mathbf{\delta_{1}}$ are sufficiently small. We define $\tilde{z}_{j}$ as the center of mass of $\tilde{C}_{1}^{j}$ and we define $\tilde{D}_{j}$ as the $\tilde{z}_{j}$-cone of $\tilde{C}_{1}^{j}$. Since the center of mass of a curve varies continuously with that curve, we can assume $\tilde{z}_{j}$ to be sufficiently close to $z_{j}$. In particular we can assume that $||z-z_{j}||<\tfrac{1}{2}$. We then get that, as parametrized disks, $||\tilde{D}_{j}-D_{j}||$ is arbitrarily small. Let $\delta_{3}>0$. For $\delta$, $\delta_{1}$, $\delta_{2}$ small enough we have
$||\tilde{C}_{1}^{j}-C_{1}^{j}||,||\tilde{D}_{j}-D_{j}||<\delta_{3}.$
@@
$\sum\limits_{i\in A_{j}}\mathrm{len}\left(\tilde{C}_{1}|_{\tilde{I}_{i}}\right)\geq$ $2\pi+2\mathrm{diam}(S_{j})-\sum_{k=1}^{c}\varphi(||\tilde{C}_{1}(\tilde{a}_{y_{k}})-\tilde{C}_{1}(\tilde{a}_{x_{k+1}})||)$
@@
where we used that $x_{k+1},y_{k}\in\alpha_{l(y_{k})}$. Let $q^{A}(l)=\mathrm{diam}(\tilde{C}_{1}(\alpha_{l}))$, $q^{B}(l)=\mathrm{diam}(\tilde{C}_{2}(\beta_{l}))$, and let $f^{A}(l)=\#\{j\mid\alpha_{l}\text{ along }C_{1}^{j}\}$, $f^{B}(l)=\#\{j\mid\beta_{l}\text{ along }C_{2}^{j}\}$. Let $n=\#\{A\text{-indices }j\}$. By summing over all $j$, we get
@@
there is $\delta_{4}>0$ such that on $[0,\delta_{4}]$ we have $F(x)\leq mx$ with equality if and only if $x=0$. In particular, since $q^{A}(l),q^{B}(l)<\delta_{1}$, for $\delta_{1}<\delta_{4}$, we have
@@
with equality if and only if $q^{A}(l)=0$ and $q^{B}(l)=0$ respectively. Thus, for $d=||\tilde{L}-L||$ small enough, we have
@@
Proof.
In the last theorem we showed that $S$ is strictly locally minimizing. It remains to be shown that $S$ is closed. Take a sequence $(L_{n})_{n\in\mathbb{N}}$ in $S$ converging to some thick link $L=(C_{1},C_{2})$. Let $\{A_{j}\},\{\alpha_{l}\},\{B_{j}\},\{\beta_{l}\}$ be the index sets and let $h^{A},h^{B}$ be the index functions belonging to $S$. For each $n\in\mathbb{N}$, consider the sets of breaking points $\{(a_{i})_{n}\},\{(b_{i})_{n}\}\subseteq S^{1}$. Since $S^{1}$ is compact, we can assume that all these sequences $((a_{i})_{n})_{n\in\mathbb{N}}$ and $((b_{i})_{n})_{n\in\mathbb{N}}$ converge, else take a subsequence of $(L_{n})_{n\in\mathbb{N}}$. We denote the limits by $a_{i}$ and $b_{i}$ respectively. Note that the $\{(a_{i})_{n}\}$ and $\{(b_{i})_{n}\}$ each have the same cyclic ordering for all $n$, hence the cyclic orderings of $\{a_{i}\}$ and $\{b_{i}\}$ are (weakly) preserved when taking limits (but a priori it seems possible that $a_{i}=a_{i+1}$ or $b_{i}=b_{i+1}$ for some $i$). We will now show that ($(\{A_{j}\},\{\alpha_{l}\},\{a_{i}\}),(\{B_{j}\},\{b_{i}\},\{\beta_{l}\})$ form a suitable conjoined circle decomposition of $L$
Claim: $\{A_{j}\}$, $\{\alpha_{l}\}$, $\{a_{i}\}$ form a circle decomposition of $C_{1}$ and $\{B_{j}\}$, $\{b_{i}\}$, $\{\beta_{l}\}$ form a circle decompositon of $C_{2}$. We show this claim for $C_{1}$, the proof for $C_{2}$ is the same. It is obvious that circles $(C_{1}^{j})_{n}$ of $(C_{1})_{n}$ converge to circles $C_{1}^{j}$ of $C_{1}$. It is left to show that
@@
It is easy to see that we have $i_{1},i_{2}\in\alpha_{l}\implies C_{1}(a_{i_{1}})=C_{1}(a_{i_{2}}).$ Now assume that $i_{1}\in\alpha_{l_{1}}$ and $i_{2}\in\alpha_{l_{2}}$ with $l_{1}\neq l_{2}$. Let $K_{j_{1}}$ and $K_{j_{2}}$ be the unit circles of $C_{2}$ with $z_{j_{1}}=C_{1}(\alpha_{l_{1}})$ and $z_{j_{2}}=C_{1}(\alpha_{l_{2}})$. Since $l_{1}\neq l_{2}$, we have that $(C_{1})_{n}(\alpha_{l_{1}})\neq(C_{1})_{n}(\alpha_{l_{2}})$ for every $n\in\mathbb{N}$. It follows that for every $n\in\mathbb{N}$ we have $(z_{j_{1}})_{n}\neq(z_{j_{2}})_{n}$ and therefore $(K_{j_{1}})_{n}\neq(K_{j_{2}})_{n}$. Since the circle decompositions of every $L_{n}$ are suitable, there is some $A$-index $j^{\prime}$ such that for every $n$ the circle $(K_{j^{\prime}})_{n}$ is non-trivially linked to $(K_{j_{1}})_{n}$, but trivially linked to $(K_{j_{2}})_{n}$. Taking limits, we get that $K_{j^{\prime}}$ is non-trivially linked to $K_{j_{1}}$ but trivially linked $K_{j_{2}}$, hence $K_{j_{1}}\neq K_{j_{2}}$. This implies $z_{j_{1}}\neq z_{j_{2}}$, since if two different circles $K_{j_{1}}$, $K_{j_{2}}$ of $C_{2}$ had the same center, it would be impossible for $C_{1}$ to pierce their common center while keeping unit distance from both. Therefore $C_{1}(a_{i_{1}})=z_{j_{1}}\neq z_{j_{2}}=C_{1}(a_{i_{2}})$. This shows that two $A$-indices $i_{1},i_{2}$ are in the same point cluster if and only if $C_{1}(a_{i_{1}})=C_{1}(a_{i_{2}})$. It is left to check that the intervals $I_{i}=[a_{i},a_{i+1}]$ all have positive length, i.e. $a_{i}\neq a_{i+1}$. If $i,i+1$ are in different point clusters, then by what we have just shown $C_{1}(a_{i})\neq C_{1}(a_{i+1})$ and hence $a_{i}\neq a_{i+1}$. If $i,i+1$ are in the same $\alpha_{l}$, then there is some $j$ with $A_{j}=\{i\}$. This means that $(C_{1})_{n}([(a_{i})_{n},(a_{i+1})_{n}])=(K_{j})_{n}$ is a unit circle for each $n$. By taking limits, $C_{1}([a_{i},a_{i+1}])=K_{j}$ is a unit circle, hence $I_{i}$ has positive length.
@@
The first condition is obviously fulfilled. We will show that the second condition is fulfilled, the third one follows analogously. Let $K_{j}$ be a circle of $C_{1}$. Let $i$ be such that $z_{j}^{n}=(C_{2})_{n}((b_{i})_{n})$ for every $n$. Then by taking limits, $z_{j}=C_{2}(b_{i})$. We omit the proof that every intersection is transverse after taking limits. The conjoined circle decomposition has the same index functions $h^{A}$ and $h^{B}$ as that of $L$. Note that taking limits preserves both thick Hopf links and thick unlinks, so we have $K_{j}\text{ linked to }K_{j^{\prime}}\iff K_{j}^{n}\text{ linked to }K_{j^{\prime}}^{n}.$ This shows that $L$ has the same linking scheme as each $L_{n}$.
@@
  • 2.
    for every $A$-index $l$: #$\#\{j\text{ }A\text{-index }\mid\alpha_{l}\text{ along }K_{j}\}<2\#(h^{A})^{-1}(l)$,
  • 3.
    for every $B$-index $l$: #$\#\{j\text{ }B\text{-index }\mid\beta_{l}\text{ along }K_{j}\}<2\#(h^{B})^{-1}(l)$,
  • 4.
@@
$\{j\text{ }A\text{-index }\mid\alpha_{l}\text{ along }C_{1}^{j}\}=\{j\text{ }A\text{-index }\mid\alpha_{l}\text{ along }(C_{1}^{j})_{n}\}.$
@@ 6.3 A Gordian unlink and other examples
A two-component Gordian unlink exists.
@@
One carefully checks that the link shown in Figure 7, parametrized as indicated in 9 has a suitable conjoined circle decomposition with linking number equal to zero. ∎
@@
For every $n\in\mathbb{N}$, there exists a $2$-component thick link $L$ with linking number $n$, which is a non-global, local minimum and a sink. In particular, there are Gordian pairs in every link homotopy class with $2$ components.
@@
Let $(C_{1},C_{2})$ be the Gordian unlink from Theorem 6.58 and with parametrization indicated as in Figure 9. If $n=0$, we can just take $L=(C_{1},C_{2})$ and we are done, so let $n\geq 1$. Let $C_{1}$ be the curve with colors yellow, green, and blue, parametrized by starting at yellow–grey breaking point of the point cluster $F$ and going along the grey segment. Let $S_{1}$ be the curve parametrizing the grey unit circle around $E$ of the sketch, parametrized with the same orientation it has as a part of $C_{1}$. Let $C_{2}$ be the curve with the colors red, purple, and orange, parametrized by starting at the red–grey breaking point of the point cluster $A$ and going along the grey segment. Let $S_{2}$ be the curve parametrizing the grey unit circle around $F$, parametrized with the same orientation it has as a part of $C_{1}$. Now consider $\tilde{C}_{1}$ defined by concatenating $k$ copies of $C_{1}$ and $n$ copies of $S_{1}$, and consider $\tilde{C}_{2}$ defined by concatenating $k$ copies of $C_{2}$ and one copy of $S_{2}$.
@@
$\mathrm{link}(\tilde{C}_{1},\tilde{C}_{2})=$ $k^{2}\mathrm{link}(C_{1},C_{2})+nk\mathrm{link}(S_{1},C_{2})+k\mathrm{link}(C_{1},S_{2})+n\mathrm{link}(S_{1},S_{2})$
$=$ $0+0+0+n\mathrm{link}(S_{1},S_{2})=n$

### v2

@@ 7.2 Circle decompositions - the details
Let $C:I\to\mathbb{R}^{3}$ be a uniformly continuous curve and $\varepsilon>0$. Then there is a $\delta>0$ such that for all $\tilde{C}$ with $\tilde{C}-C||<\delta$ and all intervals $a,b,a^{\prime},b^{\prime}\in I$ with $a<b$, $a^{\prime}<b^{\prime}$ and $|a-a^{\prime}|,|b-b^{\prime}|<\delta$, we have $\bar{C}-\bar{C}^{\prime}||<\varepsilon,$ where
@@
[…]
@@
Let $C:I\to\mathbb{R}^{3}$ be a uniformly continuous curve and $\varepsilon>0$. Then there exists a $\delta>0$ such that for every $C^{\prime}$ with $C^{\prime}-C||<\delta$ and every collection of $n$ subcurves $C_{i}:[a_{i},b_{i}]\to\mathbb{R}^{3}$, $i=1,\dots,n$ of $C$ and of $n$ subcurves $C^{\prime}_{i}:[a^{\prime}_{i},b^{\prime}_{i}]\to\mathbb{R}^{3}$ of $C^{\prime}$ with $|a_{i}-a^{\prime}_{i}|,|b_{i}-b^{\prime}_{i}|<\varepsilon$ we have
$C-C^{\prime}||<\varepsilon,$
@@
Let $\tilde{L}=\tilde{C}_{1}\cup\tilde{C}_{2}$ be a thick link with $L-\tilde{L}||=d$.
For each $A$-index $l$ and $B$-index $l$, pick closed intervals $\{H_{i}\}_{i\in\alpha_{l}}$ with $a_{i}\in H_{i}^{\circ}$, and pick closed intervals $\{H_{i}\}_{i\in\beta_{l}}$ with $b_{i}\in H_{i}^{\circ}$, such that $\{H_{i}\mid i\text{ }A\text{-index}\}$ are pairwise disjoint and such that $\{H_{i}\mid i\text{ }B\text{-index}\}$ are pairwise disjoint. Now, define sets of points $\{\tilde{a}_{i}\}_{i\in\alpha_{l}}$ such that $\tilde{a}_{i}\in H_{i}$ and $\mathrm{diam}(\tilde{C}_{1}(\{\tilde{a}_{i}\}_{i\in\alpha_{l}}))$ is minimized, and $\{\tilde{b}_{i}\}_{i\in\beta_{l}}$ with $\tilde{b}_{i}\in H_{i}$ and such that $\mathrm{diam}(\tilde{C}_{2}(\{\tilde{b}_{i}\}_{i\in\beta_{l}}))$ is minimized. Since the circle decomposition is suitable, we can assume that there are always two curves in $\{C_{1}\mid_{H_{i}}\}_{i\in\alpha_{l}}$ only intersecting in $C_{1}(\alpha_{l})$ and two curves in $\{C_{2}\mid_{H_{i}}\}_{i\in\beta_{l}}$ only intersecting in $C_{2}(\beta_{l})$. This allows us to apply Lemma 7.17 to the families of curves $\{C_{1}\mid_{H_{i}}\}_{i\in\alpha_{l}}$ and $\{C_{2}\mid_{H_{i}}\}_{i\in\beta_{l}}$. We get the following result. For any $\mathbf{\delta_{1}>0}$, there is $\mathbf{\delta>0}$ such that $\mathbf{d<\delta}$ implies
@@
$C_{1}^{j}-\tilde{C}_{1}^{j}||<\delta_{2}$
as long as $\mathbf{\delta}$ and $\mathbf{\delta_{1}}$ are sufficiently small. We define $\tilde{z}_{j}$ as the center of mass of $\tilde{C}_{1}^{j}$ and we define $\tilde{D}_{j}$ as the $\tilde{z}_{j}$-cone of $\tilde{C}_{1}^{j}$. Since the center of mass of a curve varies continuously with that curve, we can assume $\tilde{z}_{j}$ to be sufficiently close to $z_{j}$. In particular we can assume that $z-z_{j}||<\tfrac{1}{2}$. We then get that, as parametrized disks, $\tilde{D}_{j}-D_{j}||$ is arbitrarily small. Let $\delta_{3}>0$. For $\delta$, $\delta_{1}$, $\delta_{2}$ small enough we have
[…]
@@
[…]
@@
where we used that $x_{k+1},y_{k}\in\alpha_{l(y_{k})}$. Let $q^{A}(l)=\mathrm{diam}(\tilde{C}_{1}(\alpha_{l}))$, $q^{B}(l)=\mathrm{diam}(\tilde{C}_{2}(\beta_{l}))$, and $f^{A}(l)=\#\{j\mid\alpha_{l}\text{ along }C_{1}^{j}\}$, $f^{B}(l)=\#\{j\mid\beta_{l}\text{ along }C_{2}^{j}\}$. Let $n=\#\{A\text{-indices }j\}$. By summing over all $j$, we get
@@
there is $\delta_{4}>0$ such that on $0,\delta_{4}]$ we have $F(x)\leq mx$ with equality if and only if $x=0$. In particular, since $q^{A}(l),q^{B}(l)<\delta_{1}$, for $\delta_{1}<\delta_{4}$, we have
@@
[…]
@@
Proof of Theorem 7.9.
In the last theorem we showed that $S$ is strictly locally minimizing. It remains to be shown that $S$ is closed. Take a sequence $(L_{n})_{n\in\mathbb{N}}$ in $S$ converging to some thick link $L=(C_{1},C_{2})$. Let $\{A_{j}\},\{\alpha_{l}\},\{B_{j}\},\{\beta_{l}\}$ be the index sets and let $h^{A},h^{B}$ be the index functions belonging to $S$. For each $n\in\mathbb{N}$, consider the sets of breaking points $\{a_{i,n}\},\{b_{i,n}\}\subseteq S^{1}$. Since $S^{1}$ is compact, we can assume that all these sequences $(a_{i,n})_{n\in\mathbb{N}}$ and $(b_{i.n})_{n\in\mathbb{N}}$ converge, else take a subsequence of $(L_{n})_{n\in\mathbb{N}}$. We denote the limits by $a_{i}$ and $b_{i}$ respectively. Note that for each fixed $n$, the cyclic ordering of $\{a_{i,n}\}$ as well as that of $\{b_{i,n}\}$ are the same, hence the cyclic orderings of $\{a_{i}\}$ and $\{b_{i}\}$ are (weakly) preserved when taking limits (but a priori it seems possible that $a_{i}=a_{i+1}$ or $b_{i}=b_{i+1}$ for some $i$). We will now show that $(\{A_{j}\},\{\alpha_{l}\},\{a_{i}\}),(\{B_{j}\},\{b_{i}\},\{\beta_{l}\})$ form a suitable conjoined circle decomposition of $L$.
Claim: $\{A_{j}\}$, $\{\alpha_{l}\}$, $\{a_{i}\}$ form a circle decomposition of $C_{1}$ and $\{B_{j}\}$, $\{b_{i}\}$, $\{\beta_{l}\}$ form a circle decompositon of $C_{2}$. We show this claim for $C_{1}$, the proof for $C_{2}$ is the same. It is obvious that circles $C_{1,n}^{j}$ of $C_{1,n}$ converge to circles $C_{1}^{j}$ of $C_{1}$. It is left to show that
@@
It is easy to see that we have $i_{1},i_{2}\in\alpha_{l}\implies C_{1}(a_{i_{1}})=C_{1}(a_{i_{2}}).$ Now assume that $i_{1}\in\alpha_{l_{1}}$ and $i_{2}\in\alpha_{l_{2}}$ with $l_{1}\neq l_{2}$. Let $K_{j_{1}}$ and $K_{j_{2}}$ be the unit circles of $C_{2}$ with $z_{j_{1}}=C_{1}(\alpha_{l_{1}})$ and $z_{j_{2}}=C_{1}(\alpha_{l_{2}})$. Since $l_{1}\neq l_{2}$, we have that $C_{1,n}(\alpha_{l_{1}})\neq C_{1,n}(\alpha_{l_{2}})$ for every $n\in\mathbb{N}$. It follows that for every $n\in\mathbb{N}$ we have $z_{j_{1},n}\neq z_{j_{2},n}$ and therefore $K_{j_{1},n}\neq K_{j_{2},n}$. Since the circle decompositions of every link $L_{n}$ are suitable, there is some $A$-index $j^{\prime}$ such that for every $n$ the circle $K_{j^{\prime},n}$ is non-trivially linked to $K_{j_{1},n}$, but trivially linked to $K_{j_{2},n}$. Taking limits, we get that $K_{j^{\prime}}$ is non-trivially linked to $K_{j_{1}}$ but trivially linked $K_{j_{2}}$, hence $K_{j_{1}}\neq K_{j_{2}}$. This implies $z_{j_{1}}\neq z_{j_{2}}$, since if two different circles $K_{j_{1}}$, $K_{j_{2}}$ of $C_{2}$ had the same center, it would be impossible for $C_{1}$ to pierce their common center while keeping unit distance from both. Therefore $C_{1}(a_{i_{1}})=z_{j_{1}}\neq z_{j_{2}}=C_{1}(a_{i_{2}})$. This shows that two $A$-indices $i_{1},i_{2}$ are in the same point cluster if and only if $C_{1}(a_{i_{1}})=C_{1}(a_{i_{2}})$. It is left to check that the intervals $I_{i}=[a_{i},a_{i+1}]$ all have positive length, i.e. $a_{i}\neq a_{i+1}$. If $i,i+1$ are in different point clusters, then by what we have just shown $C_{1}(a_{i})\neq C_{1}(a_{i+1})$ and hence $a_{i}\neq a_{i+1}$. If $i,i+1$ are in the same $\alpha_{l}$, then there is some $j$ with $A_{j}=\{i\}$. This means that $C_{1,n}([(a_{i,n},a_{i+1,n}])=K_{j,n}$ is a unit circle for each $n$. By taking limits, $C_{1}([a_{i},a_{i+1}])=K_{j}$ is a unit circle, hence $I_{i}$ has positive length.
@@
The first condition is obviously fulfilled. We will show that the second condition is fulfilled, the third one follows analogously. Let $K_{j}$ be a circle of $C_{1}$. Let $i$ be such that $z_{j,n}=C_{2,n}(b_{i,n})$ for every $n$. Then by taking limits, $z_{j}=C_{2}(b_{i})$. That the intersection is transverse can be seen by realizing that the arclength parametrizations of $C_{2}$ is $\mathcal{C}^{1}$ and piecewise $\mathcal{C}^{2}$. This fact, together with the observation that $C_{2}$ keeps unit distance from the boundary of the unit disk $D_{j}$, means that any intersection of the arclength parametrization of $C_{2}$ with the unit disk $D_{j}$ has to be transverse.
[…]
@@
  • for every $A$-index $l$: $\{j\text{ }A\text{-index }\mid\alpha_{l}\text{ along }K_{j}\}<2\#(h^{A})^{-1}(l)$,
  • for every $B$-index $l$: $\{j\text{ }B\text{-index }\mid\beta_{l}\text{ along }K_{j}\}<2\#(h^{B})^{-1}(l)$,
@@
[…]
@@ 7.3 A Gordian unlink and other examples
7.23 Remark.
When considering $2$-component links up to link homotopy, the linking number is a complete link invariant, see [15]. This means a $2$-component link is (link homotopically) an unlink if and only if its linking number is zero.
@@
A two-component homotopically Gordian unlink exists.
@@
One carefully checks that the link shown in Figure 7, parametrized as indicated in Figure 9 has a suitable conjoined circle decomposition, and that the link has linking number equal to zero. ∎
@@
For every link homotopy class with $2$-components, there exists a thick link in that class which is a non-global, local minimum and a sink for ropelength. In particular, there are Gordian pairs in every link homotopy class with $2$ components.
@@
By Remark 7.23, it suffices to show that for every $n\in\mathbb{N}$, there exists a $2$-component thick link $L$ with linking number $n$, which is a non-global, local minimum and a sink.
Let $(C_{1},C_{2})$ be the Gordian unlink from Theorem 7.24 and with parametrization indicated as in Figure 9. If $n=0$, we can just take $L=(C_{1},C_{2})$ and we are done, so let $n\geq 1$. Let $C_{1}$ be the curve with colors yellow, green, and blue, parametrized by starting at yellow-grey breaking point of the point cluster $F$ and going along the grey segment. Let $S_{1}$ be the curve parametrizing the grey unit circle around $E$ of the sketch, parametrized with the same orientation it has as a part of $C_{1}$. Let $C_{2}$ be the curve with the colors red, purple, and orange, parametrized by starting at the red–grey breaking point of the point cluster $A$ and going along the grey segment. Let $S_{2}$ be the curve parametrizing the grey unit circle around $F$, parametrized with the same orientation it has as a part of $C_{1}$. Now consider $\tilde{C}_{1}$ defined by concatenating $k$ copies of $C_{1}$ and $n$ copies of $S_{1}$, and consider $\tilde{C}_{2}$ defined by concatenating $k$ copies of $C_{2}$ and one copy of $S_{2}$.
@@
$\mathrm{link}(\tilde{C}_{1},\tilde{C}_{2})=$ $k^{2}\cdot\mathrm{link}(C_{1},C_{2})+nk\cdot\mathrm{link}(S_{1},C_{2})$
$+$ $k\cdot\mathrm{link}(C_{1},S_{2})+n\cdot\mathrm{link}(S_{1},S_{2})$
$=$ $0+0+0+n\cdot\mathrm{link}(S_{1},S_{2})=n$
@@
Acknowledgements
The author is grateful to John Sullivan for suggesting the study of thick links related to the $4$-component configuration described in Section 6.